

********************************************************************
UARS WEEKLY STATUS REPORT - 08 DEC 2000
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible. We
continue to gather enough data to meet all the SOLSTICE science objectives
during this period of reduced tape recorder operations.
In support of the OA_SSPP_ATT patch testing, we have successfully run the
the last year's worth of data under RSS. We are encouraged by this and
analysis of these results continues.
***********
SUSIM
***********
SUSIM was turned off for the recent yaw around. Only data for
Dec. 1-2 were missed. Otherwise, SUSIM performed daily solar
scans of the sun during periods of predicted telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM/AXIS X-ray imager was operated between November 17 and November 19
in full configuration mode. Half of the imager was operated between October 16
through October 30 and November 26 through November 30. The PEM zenith
spectrometers are activate in daylight mode.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. PEM:: is now over 10 years old and its hardware
is now very difficult and costly to maintain. PEM:: is also showing signs
of slowly failing, and it is doubtful that PEM:: could be repaired if it failed.
The PEM system depends on PEM:: for instrument commanding and telemetry flow,
so it is an integral part of the PEM ground system. Current tests of telemetry
transfers through PEM2K:: have been successful. PEM Telemetry is now runing
in duplicate streams through PEM:: and PEM2K::.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in two UARS meetings on 6 December 2000.
The first meeting was called to report the status of effort to
process three-axis magnetometer data in the on-board computer and
use the on-board solution along with fine sun sensor data to
control attitude of the UARS spacecraft. At this meeting it was
reported that operation of this system will be started later this
week. The first test will use the on-board solution for about 1
to 2 orbits. Assuming that results from the first test are
acceptable, another test of about 7 hours will be scheduled.
The MPG also participated in the UARS Monthly Status Review
meeting, which was the second meeting held on 6 December.
Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now complete
to 9 February 2001.
***********
CDHF
***********
o DAAC DATA:
- Data created in the month of September, 2000 has been loaded into the
DAAC Transfer Tables for HALOE, NMC, and UKMO. DAAC has been notified.
- WINDII requested Version 11 be sent to the DAAC from start of mission to
31-Aug-1997. WINDII has completed updating the UARS_QUALITY flags.
Plans are to start loading data into the DAAC Transfer Tables. DAAC
has been notified of plans.
- Data created in the month of October, 2000 will be previewed and loaded
into the Daac Transfer Tables December, 2000.
o UARS CDHF REHOST:
- Need positive response from all PI End Users prior to Rehost transition.
- PI validation of output products are being addressed.
***********
FOT
***********
REPORTING PERIOD
This report covers 25 November 2000 (Orbit 50326, GMT Day 330) through 01
December 2000 (Orbit 50430, GMT Day 336).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 330/2336 50340 POWER OFF
PEM 331/0039 50341 AXIS 1 POWER ON
HALOE 333/2342 50385 Power Off
PEM 335/1327 50409 PEM/HEPS/ZEPS Power Off
PEM 335/1327 50409 Day Only to OFF
ACRIM 335/2344 50415 Power Off
PEM 335/2344 50415 AXIS1 OFF
SUSIM 335/2344 50415 Power Off
The beta angle ranged between the angles of 28.6 degrees to 3.7 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of 2 degrees on 3 December 2000 (DOY 337).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.6 and 1.0 degrees C respectively. Additional battery performance data
for each battery on 24 November 2000 and 1 December 2000 are provided in
Appendix D.
The clock error ranged from Ð10.6 to 6.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 17
¥ WINDII = 01
¥ PEM = 04
¥ HALOE = 01
¥ ACRIM = 01
¥ SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/221 - 2287.496937
· Transponder B: 00/305 - 2287.499087
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/335 show the observatory data loss to be 300 hours, 00
minutes, 47.608 seconds. This shows an overall increase of 1 minute and 49.008
seconds since the last report period. This is a 0.39 percent data loss, which
equals a 99.61 percent data capture for the mission.
The decrease of 1 minute and 49.008 seconds occurred over one (1) day:
· 1 minute, 49.008 seconds on Day 335 due to a Late AOS; AR00-047,TTR#14170
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT put the TAM into monitor mode on 28 November 2000 (DOY 333). Its
performance is being monitored and trended daily to verify proper operation
before it is used as an input to the update filter. So far the TAM performance
on board shows slightly higher residuals than were found on the simulator. The
FOT continues to analyze these high residuals and hopes to put the TAM into
update mode sometime next week.
MEETINGS
The FOT attended a UARS Attitude meeting on 29 November 2000. This meeting was
called to determine the criteria for putting the TAM into update mode and
review the yaw plan.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for December 31, 2000 (DOY 366). This
maneuver will be a Forward to Reverse Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 01 December 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
50337 330/1746 00-665 NBTR NOOPS D330
50337 330/1746 00-666 NBTR NOOPS D331
50340 330/2336 00-667 WINDII POWER OFF
50341 331/0039 00-668 PEM AXIS 1 POWER ON
50344 331/0544 00-666 NBTR NOOPS D331
50352 331/1749 00-666 NBTR NOOPS D331
50353 331/2013 00-669 NBTR NOOPS D332
50360 332/0735 00-669 NBTR NOOPS D332
50368 332/1939 00-670 NBTR NOOPS D333
50368 332/2016 00-671 OBC Memory Monitor Change
50375 333/0630 00-670 NBTR NOOPS D334
50380 333/1423 00-672 OBC/ACS TAM Time Values
50380 333/1423 00-673 OBC/ACS TAM Error Variance
50380 333/1423 00-674 OBC/ACS TAM to Monitor Mode
50385 333/2342 00-675 HALOE Power Off
50388 334/0405 00-676 NBTR PLAYBACK
50395 334/1505 00-677 OBC/ACS Slide MAGMON Processor
50395 334/1505 00-678 OBC/ACS Reset/Enable MAGMON Proc
50409 335/1327 00-679 PEM PEM/HEPS/ZEPS Power Off
50409 335/1327 00-680 PEM Day Only to OFF
50411 335/1613 00-681 OBC New MemMon Locations
50413 335/1945 00-682 OBC New MemMon Locations
50415 335/2344 00-683 ACRIM Power Off
50415 335/2344 00-684 PEM AXIS1 OFF
50415 335/2344 00-685 SUSIM Power Off
__________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 01 December 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 01 December 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period
__________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 01 December 2000
24 Nov 2000 - GMT Day 329 Beta = 32.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- --------- --------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/+0.0 14.2 0.000 0.0
2 +11.2/-56.0 +5.79/+4.51 +29.6/-15.2 25.1 1.030 21.1
3 +28.0/-33.6 +4.51/+3.23 +30.4/-16.0 25.0 1.030 20.7
01 Dec 2000 - GMT Day 336 Beta = 03.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- -------- --------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 14.2 0.000 0.0
2 +5.6/-100.8 +5.15/+4.51 +29.6/-13.2 25.1 1.020 21.0
3 +44.8/-11.2 +3.87/+2.92 +28.8/-13.6 25.1 1.020 21.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
__________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 01 December 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-047 FOT 50415 335/2344 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-047 FOT see list Nov 00 14170 Generic Late Acquisition
00-053 FOT 50168 319/1113 23157 Late Acquisition
00-054 FOT 50280 326/2147 23187 RESET PTP /2:39 32K
SEVNTFW00/48